Please help with factoring the following polynomials
2x to the power of 2 - 3x to the power of 2 -32x+48

Answer by nic0le116(25)

2x^2-3x^2-3x+48 Combine like terms
-1x^2-3x+48 I find it easier to work with a leading coefficient of positive one, so I would switch the signs of every number
x^2+3x-48 Then look and ask yourself, what two numbers multiply to be -48 but when these two numbers are added, they add up to be 3.
These numbers are 8 and -6.
8*(-6)=-48
8+(-6)=3
ANSWER=
(x+8)(x-6)
If you're ever unsure if your answer is correct, foil your factoring back out and see if it equals the equation before you factored.